OpenStudy (anonymous):

1.) A sheet of paper has a length of 29 cm, a width of 8 inches, and a thickness of 0.10 mm. What is the volume of the sheet of paper in cm^3? 2.) Calculate the density in kg/m^3 of an object that measures 5 cm on each side and has a mass of 320 grams.

OpenStudy (yrelhan4):

1 inch=2.5 cms.. 1mm=0.1 cm. first convert width and thickness into cms.. Volume= length*width*thickness..

OpenStudy (anonymous):

then?

OpenStudy (yrelhan4):

You can solve the first part now, right?

OpenStudy (anonymous):

Ah! Thanks :)

OpenStudy (anonymous):

How about in number 2?

OpenStudy (yrelhan4):

Yeah you know density=mass/volume.. mass=320/1000=.32kgs volume=5^3 (i am assuming the object is a cube here.. cant be anything else)..

OpenStudy (anonymous):

Again?

OpenStudy (anonymous):

density=(mass/volume). 320grams=0.32kg 5cm=0.05m so volume=(0.05)^3 so density=0.32/(0.05)^3=2560 kg/m^3